OpenStudy (anonymous):

Explain, in complete sentences, how a higher order root is different from a square root. Part 2: Give two examples

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

example of a higher order root would be cube root of 8 (cube root is of order 3) cube root of 8 = 2 * 2 * 2 If the higher order root is odd then the number can be negative eg cube root of -8 -8 = -2*-2*-2 Square roots of a positive number has 2 square root s eg 4 = 2*2 or -2 * -2 The square root of a negative number has no real roots. more examples are 5th root of 32 = 2 ( 2*2*2*2*2) 4th root of 16 = 2 (2*2*2*2) or -2 (-2*-2*-2*-2)

OpenStudy (anonymous):

thank you